OpenStudy (anonymous):

Simplify. −4r + 3s − (−2r + s) ? show work

OpenStudy (anonymous):

answer it mathematicalpro!!

OpenStudy (anonymous):

-4r+3s+2r-s =-2r+2s =-2(r-s)

OpenStudy (anonymous):

Not sure why you pulled out the negative there: 2(s-r) seems like a nicer answer.

OpenStudy (anonymous):

same difference though ;p

OpenStudy (anonymous):

both r same

OpenStudy (anonymous):

Nobody appreciates my puns..

OpenStudy (anonymous):

but i think ur answer is nice :)